What is the slope of the given supply curve? (12, S15) Price ($) (9,$9) Supply 10 Quantity


What is the slope of the given demand curve? (5, $15) Demand (13, $11) Price ($) Quantity


Use the graph to determine the units of the slope and the y-intercept 

units of slope: 

units of y-intercept:


What is the slope of the given supply curve? 


What is the slope of the given demand curve? 
Answer #1

(Question 1)

Slope = Change in y value / Change in x value = (4 - 3) / (2 - 0) = 1/2 = 0.5

y-intercept = 3

(Question 2)

Slope = Change in y value / Change in x value = (15 - 9) / (12 - 9) = 6/3 = 2

(Question 3)

Slope = Change in y value / Change in x value = (15 - 11) / (5 - 13) = 4/(-8) = -0.5
Answer #2

Screenshot (29).png

answered by: guest